# Density estimation via mixture discrepancy and moments

### v1

@@ 4 Numerical experiments
This section conducts numerical experiments to test the numerical effects of DSP-mix and MSP compared with DSP. The first four distributions are from [4]. The $L^{2}$ relative error
$\mathcal{E}_{2}[p]=\frac{\lVert p^{\text{num}}(\bm{x})-p^{\text{ref}}(\bm{x})\rVert_{2}}{\lVert p^{\text{ref}}(\bm{x})\rVert_{2}}$ (24)
is used to measure the accuracy, where $p^{\text{ref}}(\bm{x})$ is the reference solution and $p^{\text{num}}(\bm{x})$ is the numerical solution. We calculate the $L^{2}$ norm according to the partitioning $\mathcal{P}_{L}$,
$\lVert f(\bm{x})\rVert_{2}\approx\sqrt{\sum_{l=1}^{L}f(\bm{x}_{c})^{2}|\Omega_{l}|},$ (25)
where $\bm{x}_{c}$ is the center of $\Omega_{l}$. The algorithm parameters are set as follows unless otherwise stated:
  • 1.
    DSP-mix: In Eq. (8), the parameter $\theta$ determining the level of refinement of the partition is set to 0.1. In Eq. (6), the parameter $m$ specifying the number of candidate splitting points is set to 10;
  • 2.
    MSP: The tolerance parameters in Eq. (13) is $\epsilon_{1}=0.1,\epsilon_{2}=0.1,\epsilon_{3}=0.1$.
All simulations via C++ implementations run on the High-Performance Computing Platform of Peking University: 2*Intel Xeon E5-2697A-v4 (2.60GHz, 40MB Cache, 9.6GT/s QPI Speed, 16 Cores, 32 Threads), 256GB $\times$ 16.
4.1 2-D Gaussian
First we consider the 2-D Gaussian distribution [4],
$\bm{x}\sim\mathcal{N}(\boldsymbol{\mu},\boldsymbol{\Sigma})\mathbbm{1}\left\{\bm{x}\in[0,1]^{2}\right\},$ (26)
where $\boldsymbol{\mu}=(0.5,0.5)^{T}$ and $\boldsymbol{\Sigma}=[0.08,0.02;0.02,0.02]$. Figure 3 shows the numerical solutions of the DSP, DSP-mix and MSP method. The proposed DSP-mix and MSP also have the ability to adaptively partition the computational domain, similar to DSP. They feature fine grids at positions where the probability density changes significantly and sparse grids in unimportant regions. Table 1 presents the $L^{2}$ relative error $\mathcal{E}_{2}[p]$ and running time of the three methods for the 2-D Gaussian example with the number of samples $N$. The errors of all three methods decrease as the number of samples $N$ increases, and the errors of the three methods are comparable without significant order-of-magnitude gap. The running time of DSP-mix and MSP are similar and both runs approximately ten times faster than DSP.
Table 1: 2-D Gaussian: The $L^{2}$ relative error $\mathcal{E}_{2}[p]$ and running time in seconds (in parentheses) of the three methods under different sample sizes $N$.
$N$ DSP DSP-mix MSP
$1\times 10^{4}$ 0.2747 0.2135 0.3218
(0.8109) (0.0171) (0.0161)
$1\times 10^{5}$ 0.1468 0.1303 0.1313
(3.3644) (0.2074) (0.1754)
$1\times 10^{6}$ 0.0968 0.0877 0.0888
(19.6516) (2.2268) (2.0470)
(a) 2-D Gaussian, Reference.
(b) 2-D Gaussian, DSP.
(c) 2-D Gaussian, DSP-mix.
(d) 2-D Gaussian, MSP.
Figure 3: 2-D Gaussian: Numerical solutions produced by DSP-mix, MSP and DSP with $N=1\times 10^{6}$.
4.2 2-D Mixture of Gaussians
Next, we consider the 2-D mixture of Gaussians distribution [4],
$\bm{x}\sim\frac{1}{2}\sum_{i=1}^{2}\mathcal{N}\left(\boldsymbol{\mu}_{i},\boldsymbol{\Sigma}_{i}\right)\mathbbm{1}\left\{\bm{x}\in[0,1]^{2}\right\},$ (27)
where $\boldsymbol{\mu}_{1}=(0.5,0.25)^{T}$, $\boldsymbol{\mu}_{2}=(0.5,0.75)^{T},\boldsymbol{\Sigma}_{1}=\boldsymbol{\Sigma}_{2}=[0.04,0.01;0.01,0.01]$. Figure 4 presents the numerical solutions of DSP, DSP-mix and MSP. All of them possess the capability to capture the distribution structure. Table 2 shows the $L^{2}$ relative errors $\mathcal{E}_{2}[p]$ and running time of the three methods for this example. Both DSP-mix and MSP can achieve approximately a tenfold speedup compared with DSP, while keeping the same accuracy.
Table 2: 2-D Mixture of Gaussians: The $L^{2}$ relative error $\mathcal{E}_{2}[p]$ and running time in seconds (in parentheses) of the three methods under different sample sizes $N$.
$N$ DSP DSP-mix MSP
$1\times 10^{4}$ 0.2509 0.1754 0.2918
(0.8177) (0.0163) (0.0140)
$1\times 10^{5}$ 0.1067 0.0904 0.0988
(3.2465) (0.2295) (0.1795)
$1\times 10^{6}$ 0.0502 0.0349 0.0333
(19.6791) (2.4601) (2.0504)
(a) 2-D Mixture of Gaussians, Reference.
(b) 2-D Mixture of Gaussians, DSP.
(c) 2-D Mixture of Gaussians, DSP-mix.
(d) 2-D Mixture of Gaussians, MSP.
Figure 4: 2-D Mixture of Gaussians: Numerical solutions produced by DSP-mix, MSP and DSP with $N=1\times 10^{6}$.
4.3 2-D Mixture of Betas
In this numerical experiment, we test 2-D mixture of betas distribution [4],
$\bm{x}\sim\frac{1}{3}(\operatorname{beta}(2,5)\operatorname{beta}(5,2)+\operatorname{beta}(4,2)\operatorname{beta}(2,4)+\operatorname{beta}(1,3)\operatorname{beta}(3,1)),$ (28)
where $\operatorname{beta}(\alpha,\beta)$ denotes beta distribution. Figure 5 displays the numerical visualizations of this example. Similar to the previous two examples, all three methods capture the main features of the distribution. Table 3 presents the $L^{2}$ relative errors $\mathcal{E}_{2}[p]$ and the running time of the three methods for this example respectively. The computational cost of DSP-mix and MSP is only about one-tenth of that of DSP, achieving an obvious speedup while maintaining the same precision.
Table 3: 2-D Mixture of Betas: The $L^{2}$ relative error $\mathcal{E}_{2}[p]$ and running time in seconds (in parentheses) of the three methods under different sample sizes $N$.
$N$ DSP DSP-mix MSP
$1\times 10^{4}$ 0.3015 0.2469 0.3539
(0.8728) (0.0162) (0.0146)
$1\times 10^{5}$ 0.2069 0.2065 0.2066
(3.7691) (0.2699) (0.1808)
$1\times 10^{6}$ 0.1876 0.1821 0.1831
(24.7044) (2.6406) (2.0480)
(a) 2-D Mixture of Betas, Reference.
(b) 2-D Mixture of Betas, DSP.
(c) 2-D Mixture of Betas, DSP-mix.
(d) 2-D Mixture of Betas, MSP.
Figure 5: 2-D Mixture of Betas: Numerical solutions produced by DSP-mix, MSP and DSP with $N=1\times 10^{6}$.
4.4 $d$-D Mixtures of Gaussians
In this case, we test the $d$-D mixtures of Gaussians [4],

### v2

@@ 4 Numerical experiments
This section conducts numerical experiments to test the performance of tree-based density estimation. Since a thorough comparative study between DSP and KDE has been performed in [18], we mainly focus on the performance of DSP-mix and MSP compared with DSP. In Eq. (5), the parameter $m$ specifying the number of candidate splitting points is set to $64$. To evaluate the performance of DSP, DSP-mix and MSP for high-dimensional datasets, we adopt two standard error metrics as in [18]. The first is the Kullback-Leibler (KL) divergence to quantify the similarity between two probability distributions,
$\textup{KL}(p^{\text{ref}}||p^{\text{num}})=\int p^{\text{ref}}(\bm{x})\log\left(\frac{p^{\text{ref}}(\bm{x})}{p^{\text{num}}(\bm{x})}\right)\mathrm{d}\bm{x},$ (27)
which can be evaluated by the arithmetic mean of the points $(\bm{x}_{1},\bm{x}_{2},\dots,\bm{x}_{N})\sim p^{\text{ref}}$, namely,
$\textup{KL}(p^{\text{ref}}||p^{\text{num}})\approx\frac{1}{N}\sum_{i=1}^{N}\left(\log p^{\text{ref}}(\bm{x}_{i})-\log p^{\text{num}}(\bm{x}_{i})\right).$ (28)
The second metric is the Hellinger distance,
${H}(p^{\text{ref}},p^{\text{num}})=\frac{1}{2}\int\left(\sqrt{p^{\text{ref}}(\bm{x})}-\sqrt{p^{\text{num}}(\bm{x})}\right)^{2}\textup{d}\bm{x}=1-\int\sqrt{\frac{p^{\text{num}}(\bm{x})}{p^{\text{ref}}(\bm{x})}}p^{\text{ref}}(\bm{x})\textup{d}\bm{x},$ (29)
which can be approximated by
$\hat{H}(p^{\text{ref}},p^{\text{num}})\approx\hat{H}^{2}=1-\frac{1}{N}\sum_{i=1}^{N}\sqrt{\frac{p^{\text{num}}(\bm{x}_{i})}{p^{\text{ref}}(\bm{x}_{i})}}.$ (30)
For each setting of parameters, we run 10 simulations from independent sampling point sets and measure the standard derivation
$\textup{SD}=\sqrt{\frac{1}{10}\sum_{i=1}^{10}\left(\textup{KL}^{(i)}\right)^{2}-\left(\frac{1}{10}\sum_{i=1}^{10}\textup{KL}^{(i)}\right)^{2}},$ (31)
or
$\textup{SD}=\sqrt{\frac{1}{10}\sum_{i=1}^{10}\left(\hat{H}^{(i)}\right)^{2}-\left(\frac{1}{10}\sum_{i=1}^{10}\hat{H}^{(i)}\right)^{2}}.$ (32)
In practice, we record the maximal values of the standard derivations produced by DSP, DSP-mix and MSP (maxSD for short) to verify their robustness.
All simulations via C++ implementations run on the platform: AMD Ryzen 9950x (16 cores, 32 threads) with 256GB Memory (5600Mhz). The average computational time in different scenarios is recorded in Tables 2 and 4 (Beta mixtures), Tables 6 and 8(Gaussian mixtures) as well as Tables 10 and 12 (Cauchy mixtures). When the sample size is relatively small ($N\leq 10^{6}$), the computational costs for three methods are comparable. However, for large sample size ($N\geq 10^{7}$), both DSP-mix and MSP are evidently faster than DSP, and MSP even gains an increase in speed by a factor of two to twenty. Moreover, it is observed that the computational costs of both DSP-mix and MSP scale nearly linearly with respect to the sample size.
4.1 $d$-D Beta mixtures
First, we examine the $d$-D mixtures of betas,
Figure 5 visualizes this example for $d=2$. All three methods capture the main features ( three peaks) of the distribution.
(a) KL divergence ($d=2$) (left: DSP, middle: DSP-mix, right: MSP).
(b) Hellinger distance ($d=2$) (left: DSP, middle: DSP-mix, right: MSP).
Figure 4: 22-D Beta mixtures: The KL divergence and Hellinger distance under different NN and $\theta$.
In Figure 4, it is shown that the errors of the three methods are comparable, all decreasing as sample size $N$ increases. For fixed $N$, when the parameter $\theta$ decreases within a certain range, we can obtain finer partitioning and more accurate density estimation results. However, when $\theta$ goes too small, the error will increase instead, which is an indicator of the overfitting problem. Fortunately, the overfitting can be alleviated using a larger sample size.
(a) True density and samples.
(b) Adaptive partition (left: DSP, middle: DSP-mix, right: MSP).
(c) Density estimation (left: DSP, middle: DSP-mix, right: MSP).
Figure 5: 2-D Beta mixtures: Adaptive partitions and density estimators produced by DSP-mix, MSP and DSP with $N=1\times 10^{4}$ and $\theta=0.2$.
For high-dimensional tests with $d=15,20,30$, Figure 6 presents the convergence of error metrics with respect to sample size $N=10^{4},10^{5},10^{6},10^{7},10^{8}$, with the raw data collected in Tables 1 and 3. Figure 7 plots the partition level under different $d$, NN, with $\theta$ fixed to be $0.002$ with the raw data collected in Tables 2 and 4. From the above results, we make several observations.
  1. (1)
    All three methods are very robust as the standard derivations are much smaller than the mean errors.
  2. (2)
    The accuracy of all methods can be improved by increasing sample size $N$, which validates the theoretical error bound.
  3. (3)
    The accuracy of tree-based density estimation diminishes as dimension $d$ increases. This is because high-dimensional probability density tends to be concentrated in a very localized region (the typical set).
  4. (4)
    The performance of DSP and MSP is comparable, while MSP is significantly faster. DSP-mix works well for $d\leq 6$, but becomes less accurate as dimension $d$ increases. The reason is that the mixture discrepancy is a relaxation of the star discrepancy, so that the threshold $D^{\textup{mix}}\leq\frac{\theta\sqrt{N}}{n}$ is more readily to attain. As a result, the partition level decreases, which is confirmed in Figure 7.
  5. (5)
    The partition level $L$ increases as a smaller $\theta$ is chosen. When $N$ is large, it is seen that $L\sim\theta^{-1}$. When $\theta$ is too small, the partition ceases to be split further.
  6. (6)
    When $L$ is too large, the overpartitioning of space dramatically increases the computational cost but hampers the efficiency, due to the overfitting problem.
  7. (7)
    The overfitting problem can be alleviated by increasing sample size $N$.
(a) KL divergence (left: DSP, middle: DSP-mix, right: MSP).
(b) Hellinger distance (left: DSP, middle: DSP-mix, right: MSP).
(c) Partition level (left: DSP, middle: DSP-mix, right: MSP).
Figure 6: $d$-D Beta mixtures: Convergence of DSP, DSP-mix and MSP and the partition level $L$ with respect to NN for $d=15,20,30$.
(a) Partition level ($N=10^{4}$) (left: DSP, middle: DSP-mix, right: MSP).
(b) Partition level ($N=10^{5}$) (left: DSP, middle: DSP-mix, right: MSP).
(c) Partition level ($N=10^{6}$) (left: DSP, middle: DSP-mix, right: MSP).
(d) Partition level ($N=10^{7}$) (left: DSP, middle: DSP-mix, right: MSP).
Figure 7: $d$-D Beta mixtures: The partition level $L$ under different $\theta$, $N$ and $d$. It is seen that $L\sim\theta^{-1}$ when $N$ is sufficiently large.
Table 1: $d$-D Beta mixtures: The KL divergence and Hellinger distance under different $d$ and $\theta$. The sample size is fixed to be $N=10^{7}$.
$\theta$ $0.002$ $0.02$ $0.05$ $0.1$ $0.2$
$d$ Method KL $\hat{H}^{2}$ KL $\hat{H}^{2}$ KL $\hat{H}^{2}$ KL $\hat{H}^{2}$ KL $\hat{H}^{2}$
2 DSP -0.0007 0.0011 0.0013 0.0008 0.0019 0.0008 0.0028 0.0010 0.0042 0.0013
DSP-mix 0.0001 0.0010 0.0017 0.0009 0.0029 0.0011 0.0047 0.0015 0.0087 0.0026
MSP 0.0001 0.0010 0.0011 0.0008 0.0016 0.0008 0.0023 0.0009 0.0033 0.0011
(maxSD) 0.0002 0.0001 0.0001 0.0000 0.0001 0.0001 0.0002 0.0001 0.0002 0.0001
4 DSP -0.2145 0.0542 0.0133 0.0075 0.0320 0.0105 0.0485 0.0140 0.0695 0.0192
DSP-mix -0.0546 0.0168 0.0246 0.0101 0.0501 0.0151 0.0802 0.0220 0.1201 0.0319
MSP -0.3167 0.0784 -0.0043 0.0076 0.0155 0.0078 0.0274 0.0096 0.0404 0.0123
(maxSD) 0.0012 0.0004 0.0002 0.0001 0.0006 0.0002 0.0008 0.0003 0.0012 0.0005
6 DSP -0.3601 0.0976 0.0914 0.0314 0.1490 0.0424 0.1978 0.0535 0.2531 0.0669
DSP-mix -0.0181 0.0328 0.1506 0.0441 0.2303 0.0614 0.3240 0.0822 0.4245 0.1053
MSP -0.4512 0.1166 0.0074 0.0291 0.0710 0.0306 0.1116 0.0362 0.1556 0.0445
(maxSD) 0.0003 0.0005 0.0010 0.0003 0.0015 0.0004 0.0028 0.0009 0.0052 0.0013
8 DSP -0.4151 0.1391 0.2459 0.0756 0.3636 0.0972 0.4529 0.1171 0.5552 0.1406
DSP-mix 0.0943 0.0677 0.3946 0.1043 0.5627 0.1387 0.7416 0.1738 0.8929 0.2052
MSP -0.4521 0.1458 0.0996 0.0670 0.2242 0.0752 0.1116 0.0362 0.3953 0.1043
(maxSD) 0.0008 0.0006 0.0029 0.0008 0.0058 0.0022 0.0130 0.0029 0.0126 0.0034
10 DSP -0.3763 0.1862 0.4781 0.1359 0.6679 0.1687 0.8025 0.1957 0.9590 0.2284
DSP-mix 0.2906 0.1210 0.7491 0.1838 1.0370 0.2353 1.2924 0.2779 1.4916 0.3158
MSP -0.3842 0.1875 0.2912 0.1211 0.4852 0.1386 0.6142 0.1582 0.7492 0.1838
(maxSD) 0.0015 0.0010 0.0048 0.0013 0.0126 0.0027 0.0175 0.0027 0.0165 0.0028
15 DSP 0.1017 0.3404 1.4157 0.3311 1.8129 0.3810 2.0849 0.4222 2.4313 0.4713
DSP-mix 1.1749 0.3092 2.0996 0.4180 2.7920 0.4909 3.1733 0.5313 3.5141 0.5784
MSP 0.1017 0.3406 1.1749 0.3097 1.5681 0.3443 1.8150 0.3765 2.0996 0.4179
(maxSD) 0.0099 0.0026 0.0220 0.0061 0.0326 0.0079 0.0286 0.0090 0.0345 0.0092
20 DSP 1.0627 0.5251 2.8794 0.5400 3.4996 0.5934 3.9360 0.6369 4.5112 0.6807
DSP-mix 2.6627 0.5242 4.1647 0.6455 5.2001 0.6993 5.7049 0.7391 6.2575 0.7851
MSP 1.0627 0.5257 2.6627 0.5237 3.2329 0.5605 3.6396 0.6003 4.1647 0.6452
(maxSD) 0.0195 0.0027 0.0213 0.0032 0.0205 0.0026 0.0269 0.0033 0.0295 0.0032
30 DSP 4.2883 0.8201 7.2776 0.8408 8.2751 0.8690 8.9656 0.8886 10.1624 0.9139
DSP-mix 7.1310 0.8337 9.8885 0.8990 11.3890 0.9258 12.1404 0.9391 12.9147 0.9523
MSP 4.2883 0.8195 7.1310 0.8342 8.0927 0.8590 8.7329 0.8761 9.8885 0.8989
(maxSD) 0.0382 0.0027 0.0289 0.0017 0.0417 0.0017 0.0296 0.0018 0.0252 0.0012
Table 2: $d$-D Beta mixtures: The partition level $L$ and computational time (in seconds) under different $d$ and $\theta$. The sample size is fixed to be $N=10^{7}$.
$\theta$ $0.002$ $0.02$ $0.05$ $0.1$ $0.2$
$d$ Method $L$ time $L$ time $L$ time $L$ time $L$ time
2 DSP 11601 19.34 4115 20.62 2969 27.54 2262 39.84 1634 66.47
DSP-mix 8874 15.35 4059 15.23 2915 15.75 2221 19.80 1590 31.96
MSP 8874 14.08 4342 14.31 3057 15.00 2334 14.74 1702 14.61
4 DSP 850435 947.49 28431 291.64 11696 292.27 6119 306.78 3275 329.77
DSP-mix 255326 41.38 25936 43.71 10381 62.42 5213 96.96 2619 165.91
MSP 1287088 49.38 65181 28.12 25332 27.64 14790 26.30 9079 25.90
6 DSP 1373682 1146.49 41757 430.30 17211 392.71 9272 393.10 5246 420.23
DSP-mix 270966 56.15 27402 62.56 11188 89.92 $56$ 140.38 2935 239.13
MSP 1731318 71.70 188819 42.25 76807 42.46 41359 40.14 22581 38.46
8 DSP 1707927 970.22 65024 595.31 26072 514.57 14138 489.84 8032 519.53
DSP-mix 285207 68.54 29847 80.34 12204 117.71 6260 179.82 3359 303.59
MSP 1845215 91.44 269574 59.93 108963 55.74 41359 40.14 29197 50.70
10 DSP 1921213 846.16 93104 751.74 36728 632.04 19901 585.76 11259 605.56
DSP-mix 303408 83.33 32637 96.91 13314 142.84 6924 216.97 3811 362.10
MSP 1949089 110.36 301774 72.19 122332 69.06 63136 65.12 32591 62.86
15 DSP 2235855 698.09 174225 1090.08 67961 875.88 36638 809.21 20170 859.77
DSP-mix 356114 121.08 40629 140.78 15801 204.62 8697 300.24 4900 507.12
MSP 2235891 166.61 356111 108.13 147379 100.36 77649 94.37 40629 91.47
20 DSP 2565816 723.45 268238 1436.26 105010 1174.55 56290 1083.46 30392 2503.68
DSP-mix 408780 160.17 47167 184.88 18688 260.86 10325 385.46 5456 652.46
MSP 2565816 226.91 408780 143.17 175550 134.04 91643 125.57 47167 121.13
30 DSP 3277027 898.88 425546 2079.65 169021 1626.90 95058 1825.46 45432 2000.92
DSP-mix 506018 236.90 57427 269.54 21036 379.71 11243 548.21 6627 914.78
MSP 3277027 359.31 506018 218.18 208819 196.87 116137 184.91 57427 178.54
Table 3: $d$-D Beta mixtures: The KL divergence and Hellinger distance under different $d$ and $N$. The parameter $\theta$ is fixed to be $0.002$.
NN $\times 10^{4} 1×1$$\times 10^{5}$ $\times 10^{6} 1×1$$\times 10^{7}$ $1\times 10^{8}$
$d$ Method KL $\hat{H}^{2}$ KL $\hat{H}^{2}$ KL $\hat{H}^{2}$ KL $\hat{H}^{2}$ KL $\hat{H}^{2}$
15 DSP 2.0391 0.6561 1.2169 0.5513 1.4069 0.4060 0.1017 0.3404 0.4820 0.2105
DSP-mix 2.0391 0.6554 1.2169 0.5516 1.1811 0.4085 1.1749 0.3092 1.0136 0.2461
MSP 2.0391 0.6558 1.2169 0.5508 0.5660 0.4338 0.1017 0.3406 0.2728 0.2262
20 DSP 4.0528 0.8026 2.9441 0.7384 1.9095 0.6322 1.0627 0.5251 1.2661 0.3904
DSP-mix 4.0528 0.8027 2.9441 0.7382 2.7194 0.6178 2.6627 0.5242 2.2509 0.4465
MSP 4.0528 0.8030 2.9441 0.7382 1.9095 0.6320 1.0627 0.5257 1.0996 0.3998
30 DSP 9.0334 0.9385 7.6243 0.9253 5.8924 0.8828 4.2883 0.8201 3.6955 0.7627
DSP-mix 9.0334 0.9377 7.6243 0.9254 7.1100 0.8798 7.1310 0.8337 6.2093 0.7810
MSP 9.0334 0.9380 7.6243 0.9253 5.8924 0.8828 4.2883 0.8195 4.0776 0.7293
Table 4: $d$-D Beta mixtures: The partition level $L$ and computational time (in seconds) under different $d$ and $N$. The parameter $\theta$ is fixed to be $0.002$.
NN $\times 10^{4} 1×1$$\times 10^{5}$ $\times 10^{6} 1×1$$\times 10^{7}$ $1\times 10^{8}$
$d$ Method $L$ time $L$ time $L$ time $L$ time $L$ time
15 DSP 3289 0.14 28088 1.37 83373 161.28 2235855 698.09 5850719 29904.56
DSP-mix 3289 0.15 28088 1.47 122187 13.37 356114 121.08 1058493 1335.90
MSP 3289 0.15 28088 1.35 242887 15.29 2235891 166.61 10274742 1434.56
20 DSP 3782 0.20 32780 1.95 280051 20.97 2565816 723.45 8624426 37177.42
DSP-mix 3782 0.21 32780 2.09 143972 18.12 408780 160.17 1249503 1720.63
MSP 3782 0.21 32780 1.98 280051 21.12 2565816 226.91 11640399 1854.46
30 DSP $46$ 0.34 40468 3.27 352382 34.25 3277027 898.88 11265426 38853.64
DSP-mix $46$ 0.35 40468 3.61 183723 28.50 506018 236.90 1644069 2407.87
MSP $46$ 0.35 40468 3.44 352382 34.30 3277027 359.31 14813816 2909.42
4.2 $d$-D Gaussian mixtures
Now consider the $d$-D Gaussian mixtures [18],